# Chaos and spatial prethermalization in driven-dissipative bosonic chains

### v2

@@ II.3 Two-stage thermalization in space
Let us stress that the above steady-state impurity modeling is not unique. Notably, we also tested a generalized version of the Scully-Lamb model [65] defined by $\hat{L}_{\rm s}=0$ and a modified $\hat{L}_{\uparrow}={\hat{a}^{\dagger}(\gamma^{\uparrow}-\mathcal{S}\,\hat{a}\hat{a}^{\dagger})}/{\sqrt{\gamma^{\uparrow}}}$, where $\mathcal{S}\geq 0$ is a photon saturation rate. In the context of lasing, this model can be derived as an effective theory for the optical degree of freedom when the (inverted) atomic population modeled by two-level systems has been integrated out [66, 67]. This ansatz proved to be equally successful in capturing $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$, yielding comparable values of the effective parameters $\gamma^{\uparrow}$ and $\gamma^{\downarrow}$.
II.4 Prethermal phase
We now use the above mapping to the driven-dissipative impurity ansatz defined in Eqs. (8-9) as a local thermometer along the chain. The effective inverse temperature $\beta^{\rm eff}_{\ell}$ and chemical potential $\mu^{\rm eff}_{\ell}$ at site $\ell$ can be related to the single-photon rates via the detailed-balance condition (we set $k_{\rm B}=1$)
$\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}:=\log(\gamma^{\uparrow}_{\ell}\big{/}\gamma^{\downarrow}_{\ell})\,,$ (10)
where $\gamma^{\downarrow}_{\ell}$ and $\gamma^{\uparrow}_{\ell}$ are the parameters of the impurity model determined at site $\ell$. We absorbed all the quadratic contributions to $\hat{H}_{\rm imp}$, namely $\omega_{0}-U/2$, into a redefinition of the effective chemical potential.
In Fig. 5 (a), we plot the product $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}$ as a function of $\ell$ and we identify three distinct spatial domains:
@@ II.4 Prethermal phase
    Nonsymmetric nonthermal domain; as discussed above, the first few sites are not captured by the impurity ansatz because of the proximity to the U(1)-breaking drive. This breaking of U(1) is visible in the asymmetry of local Wigner functions represented in Fig. 5 (c). Here, the quantity $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}$ is ill-defined, as indicated by the shaded region in the figure.
@@
    U(1)-symmetric prethermal domain; in this intermediate region of the chain, local quantities are well captured by the impurity ansatz with a positive chemical potential, $\mu^{\rm eff}_{\ell}>0$, and a finite 2-photon decay rate $\gamma^{\rm s}$. The relatively constant number of photons reported in Fig. 4 (a), along with the large ring-shaped Wigner functions in Fig. 5 (d), can be interpreted as the result of the competition between a large chemical potential that strives to add photons and the Kerr non-linearity that acts as a saturation mechanism. However, this state should not be identified to lasing since here, unlike lasers, the temporal phase coherence is short-lived, indicating that the U(1) symmetry breaking characteristic of lasing has not occurred (see the Supplementary Information). Instead, the phase evolves diffusively [68], restoring the U(1) symmetry of the chain Hamiltonian.
@@
    U(1)-symmetric thermal domain; the right side of the chain is captured by the impurity ansatz with a negative chemical potential, $\mu^{\rm eff}_{\ell}\leq 0$. The corresponding Wigner functions, illustrated in Fig. 5 (e), display bell-shaped envelops that can be interpreted as the outcome of a competition between the chemical potential depleting the photon population towards the vacuum, and the thermal fluctuations that sustain a residual population of weakly-interacting photons.
To test the robustness of this hydrodynamic description, we repeated the analysis above using the generalized Scully-Lamb ansatz instead of the 2-photon ansatz, as well as directly fitting the local Wigner functions with those of Gibbs states
$\hat{\rho}^{\rm eq}_{\ell}=\frac{\exp[-\beta_{\ell}^{\rm eff}(\hat{H}_{\rm imp}-\mu_{\ell}\,\hat{a}^{\dagger}a)]}{\operatorname{Tr}\left(\exp[-\beta_{\ell}^{\rm eff}(\hat{H}_{\rm imp}-\mu_{\ell}\,\hat{a}^{\dagger}a)]\right)}\,.$ (11)
A detailed comparison between these three thermometers can be found in the Supplementary Information. As reported in Fig. 5 (a), we obtained an excellent qualitative agreement between these approaches and even quantitative agreement in the thermal phase.
@@
Remarkably, we argue in Fig. 5 (b) that the sizes of both the nonsymmetric and thermal domains are limited to sub-dominant portions of the chain, whereas the size of the prethermal domain increases linearly with the overall chain length $L$. This suggests that the latter may be an extensive thermodynamic phase that occurs ahead of the eventual complete thermalization at the rightmost portion of the chain.
@@ III Discussion
The NESS of the boundary-driven dissipative chain of nonlinear oscillators revealed an extensive and chaotic region of the chain that can be effectively described hydrodynamically in terms of local equilibria with a large, emergent, chemical potential. Drawing an analogy with the long-lived prethermal phases observed in the temporal relaxation of many-body systems following a quench, this can be interpreted as a prethermal phase of matter that builds up in space rather than in time.
An immediate question concerns how this prethermal phase influences the transport properties of the driven-dissipative Bose-Hubbard chain [31, 32, 69, 70, 71]. Akin to the thermodynamic description obtained with the single-site ansatz, transport properties could be captured and analyzed using multi-site impurity models. It remains to be determined whether the length of the prethermal domain scales thermodynamically, or instead saturates at a large but finite value in very large chains. The former scenario seems plausible if the slow relaxation of the steady-state photon density across the chain is primarily governed by the distant driven-dissipative conditions at the two boundaries of the chain. However, the latter scenario is supported by the analogy to standard prethermalization mechanisms, where the timescale for escaping the prethermal state typically follows an exponential dependence, $\tau\sim\exp(1/\delta)$, with $\delta$ quantifying the strength of the integrability-breaking perturbation—here, characterized by $U/J$. Resolving this question, along with assessing the relevance of next-to-leading-order quantum fluctuations that were neglected so far, requires studying system sizes $L\gg\exp(J/U)$, well beyond current experimental capabilities. This challenge will necessitate significantly larger numerical simulations or the development of new methodologies that extend beyond the TWA to effectively handle intermediate Kerr nonlinearities while still incorporating quantum fluctuations.
More broadly, our proposed mechanism indicates that the emergence of a prethermal chaotic phase hinges on three essential conditions: first, an interacting bulk with at least one conserved charge and a sufficiently large and complex local Hilbert space that can facilitate the relaxation of different degrees of freedom across well-separated length scales; second, a local nonequilibrium drive that explicitly breaks the symmetry associated with the conservation law; and finally, weakly-symmetric dissipation channels positioned far from the drive, ensuring nontrivial steady states and a prolonged hydrodynamic relaxation of the conserved charge. These minimal ingredients, along with the semiclassical methodology we introduced, lay the foundation for identifying prethermal chaos in a wide range of boundary-driven systems.
@@ IV.1 Stochastic semiclassical description
The TWA is a semiclassical approximation of the quantum many-body dynamics that accounts for leading-order quantum fluctuations. It relies on a phase-space representation of the system’s density matrix $\hat{\rho}$ in terms of the Wigner function $W(\alpha_{1},\alpha_{1}^{*},...,\alpha_{L},\alpha_{L}^{*})$, where $\alpha_{\ell}$ and $\alpha_{\ell}^{*}$ for $\ell=1,...,L$ are the complex amplitudes associated with the local coherent states. In this framework, the Lindblad master equation on the operator $\hat{\rho}$ in Eq. (1) is mapped to a partial differential equation on $W$. Notably, two-body interactions in the Hamiltonian yield contributions up to third order derivatives of the type $\alpha_{\ell}\,\partial^{3}W/\partial\alpha^{*}_{\ell}\partial^{2}\alpha_{\ell}$. The approximation consists in discarding third and higher-order derivatives, reducing the dynamics to a Fokker-Plank equation where $W$ can be interpreted as a probability distribution of the phase-space variables. The latter equation can then be unraveled into a set of $L$ coupled stochastic differential equations on the complex amplitudes $\alpha_{\ell}$ given by Eq. (II.1). In practice, we compute the solutions of these Langevin-like equations, the so-called Wigner trajectories, by means of numerical solvers specific to stochastic differential equations. Observables are computed by averaging over a large number of trajectories generated by different realizations of the quantum noise. The dictionary between the original Lindblad master equation framework and the Wigner-trajectory implementation of the TWA framework reads, notably,
@@
where delimited-⟨$\langle$..$\rangle$ … ⟩ in the observable column is the standard quantum expectation value and, in the Wigner trajectories column, denotes the average with respect to Wigner trajectories. The local Wigner function $W_{\ell}(t;\alpha,\alpha^{*})$ is a phase-space representation of the reduced density matrix at site $\ell$, $\hat{\rho}_{\ell}(t):=\textrm{Tr}_{k\neq\ell}\,\hat{\rho}(t)$. It can be simply reconstructed by generating the histogram of $\alpha_{\ell}(t)$ in the complex plane when sampling over Wigner trajectories. When the NESS is reached, single-time observables converge to constant values and $W_{\ell}(t;\alpha,\alpha^{*})\to W^{\rm ss}_{\ell}(\alpha,\alpha^{*})$. There, the statistics can be improved by also sampling the trajectories in time, significantly reducing the computational overhead.
Figure 6: Scrambling of information in the NESS. Space-time evolution of the steady-state phase OTOC $D_{1,\ell}^{\rm ss}(\tau)$, defined in Eq. (14) and computed via Eq. (5), across an $L=80$ chain in the chaotic regime. The white dashed line indicates ballistic spreading of information at the butterfly velocity $v=2J$. The results are obtained upon averaging over $5\times 10^{3}$ independent Wigner trajectories. The drive strength has been fixed to $F=7.5$ and the other parameters are set as in Fig. 2.
IV.2 Phase OTOC
Here, we justify the expression of the phase OTOC $D_{k,\ell}(t,\tau)$ given in Eq. (5) by deriving it as the semiclassical limit of the following quantum OTOC,
$C_{k,\ell}(t,\tau)\!:=\!\frac{1}{2}\textrm{Tr}\Big{(}[\hat{W}_{\ell}(t\!+\!\tau),\hat{V}_{k}(t)]^{\dagger}[\hat{W}_{\ell}(t\!+\!\tau),\hat{V}_{k}(t)]\hat{\rho}(t)\Big{)}.$ (12)

### v3

@@ II.3.2 Hydrodynamic description
We stress that the above steady-state impurity modeling is not unique. In particular, we also tested a generalized version of the Scully-Lamb model [82, 83] defined by $\hat{L}_{\rm s}=0$ and a modified $\hat{L}_{\uparrow}={\hat{a}^{\dagger}(\gamma^{\uparrow}-\mathcal{S}\,\hat{a}\hat{a}^{\dagger})}/{\sqrt{\gamma^{\uparrow}}}$, where $\mathcal{S}\geq 0$ is a photon saturation rate. In the context of lasing, this model can be derived as an effective theory for the optical degree of freedom when the (inverted) atomic population modeled by two-level systems has been integrated out [84, 85]. This ansatz proved to be equally successful in capturing both the statics and the dynamical fluctuations yielding, notably, comparable values of the effective parameters $\gamma^{\uparrow}_{\ell}$ and $\gamma^{\downarrow}_{\ell}$.
A detailed comparison between these three thermometers, and temperature computed from the equipartition theorem whenever the latter is applicable, can be found in the Supplementary Information. As reported in Fig. 4 (d), we obtained an excellent agreement between these multiple approaches.
II.3.3 Nonthermal, prethermal, and thermal domains
The above generalized thermometers provide the spatial profiles of the local temperature and chemical potential, $T_{\ell}$ and $\mu_{\ell}$, respectively, which we use to analyze the thermodynamics along the chain, see Figs. 4 (a) and (b). In Fig. 4 (c), we present the dimensionless ratio $T_{\ell}/\mu_{\ell}$. As a complementary observable, we also monitor the entropy density $S_{\ell}:=-\textrm{Tr}[\hat{\rho}_{\ell}\log\hat{\rho}_{\ell}]$, see Fig. 4 (c), that can be computed independently, i.e. without relying on the thermodynamic ansätze.
Depending on the strength of the drive, we distinguish two regimes. At relatively weak drives ($F\lesssim 6.5$), $\mu_{\ell}$ remains negative throughout the chain and $S_{\ell}$ monotonously decreases along the chain. This is consistent with the monotonously decaying photon density profiles shown in Fig. 3 (a): negative chemical potentials lead to photon depletion toward a vacuum state dressed by thermal fluctuations. At intermediate drives ($F\gtrsim 6.5$), $\mu_{\ell}$ becomes positive and approximately constant across most of the chain, before dropping abruptly to negative values near the non-driven right boundary. This correlates with the photon density profiles shown in Fig. 3 (a): large positive chemical potentials sustain a significant photon population limited only by the Kerr nonlinearity. Concurrently, $T_{\ell}$ increases across most of the chain, reaching very high values before dropping sharply near the right end. Although the low temperatures observed at both boundaries are consistent with the proximity of the zero-temperature dissipative baths, the magnitude and position of the temperature maximum may at first seem to challenge the Clausius principle. This anomalous profile is corroborated by the entropy density profile, $S_{\ell}$, which exhibits an accumulation of entropy precisely where $T_{\ell}$ attains its peak. This counterintuitive behavior motivates our designation of this anomalous regime as “prethermal”. Similar anomalous heating profiles have previously been reported in the context of the discrete nonlinear Schrödinger equation subjected to incoherent thermal reservoirs at the two ends of a one-dimensional chain [58]. This effect has been attributed to the coupling between heat and particle transport. Specifically, it arises from the coexistence of two conserved quantities in the bulk, photon number and energy, whose associated Noether currents, $\bm{j}_{n}$ and $\bm{j}_{\epsilon}$, are coupled at the level of linear response theory.
Figure 5: Onset of the RNW Regime. Spatiotemporal evolution of (a) the photon density $n_{\ell}$, and (b) the average phase $\langle\varphi_{\ell}\rangle$ as functions of time $t$ and site index $\ell$ in a $L=100$ chain in the RNW regime. The hatched region in (b) indicates the transient chaotic regime in which the $\mathbb{U}(1)$ symmetry is restored and the average phase becomes ill-defined. Results are obtained upon averaging over $N_{\rm traj}=5\times 10^{3}$ independent Wigner trajectories. The drive amplitude is fixed to $F=12.5$. The other parameters are set as in Fig. 2.
Figure 6: RNW regime: phase decoherence from quantum fluctuations. (a) Phase fluctuations along the chain, quantified by the circular variance $\Delta\varphi_{\ell}:=1-|\langle{\rm e}^{{\rm i}\varphi_{\ell}}\rangle|$, plotted as a function of site index $\ell$ for a chain of length $L=100$ in the RNW regime. The solid blue line corresponds to the TWA results, while the dashed black line shows the classical Gross-Pitaevskii solution. (b–e) Local Wigner functions $W_{\ell}(\alpha,\alpha^{*})$ at representative sites throughout the chain. We show the normalized distribution $\tilde{W}(\alpha,\alpha^{*}):=W_{\ell}(\alpha,\alpha^{*})/\max[W_{\ell}(\alpha,\alpha^{*})]$. Yellow markers denote the classical Gross-Pitaevskii solutions. (f) Same as in panel (a), but for the first-order coherence function $|g^{(1)}_{k\ell}|$ with $k=1$. Results are computed by averaging over $N_{\rm traj}=5\times 10^{3}$ independent Wigner trajectories and over a time window $\Delta\tau=10^{3}$ once the steady state is reached. The drive amplitude is fixed to $F=12.5$. The other parameters are set as in Fig. 2.
Altogether, we identify three distinct spatial domains in the chaotic chain, from the left to the right:
@@ II.3.3 Nonthermal, prethermal, and thermal domains
    Nonsymmetric nonthermal domain; The first few sites are not captured by the different hydrodynamic ansätze because of the proximity to the $\mathbb{U}(1)$-breaking drive. This breaking of $\mathbb{U}(1)$ symmetry is visible in the asymmetry of local Wigner functions represented in Fig. 3 (d). Here, the quantities $T_{\ell}$ and $\mu_{\ell}$ are ill-defined, as indicated by the hatched region in the figure.
@@
    $\mathbb{U}(1)$-symmetric prethermal domain; In this intermediate region of the chain, only realized under strong enough driving, local quantities are well captured by hydrodynamics with a positive effective chemical potential, $\mu_{\ell}>0$, and a finite 2-photon decay rate $\gamma^{\rm s}$. The relatively constant number of photons reported in Fig. 3 (a), along with the large ring-shaped Wigner functions in Fig. 3 (e), can be interpreted as the result of the competition between a large chemical potential that strives to add photons and the Kerr non-linearity that acts as a saturation mechanism. This state should not be interpreted as lasing: unlike standard lasers, the temporal phase coherence here is short-lived, and the $\mathbb{U}(1)$ symmetry remains unbroken (see Supplementary Information). Instead, the phase undergoes diffusive dynamics [86], which restores the symmetry of the chain Hamiltonian. From a thermodynamic perspective, this domain exhibits anomalous heating, characterized by temperature $T_{\ell}$ and entropy $S_{\ell}$ that increase with distance from the coherent drive. This originates from the fact that the energy current is predominantly transported by a flux of photons with large chemical potential, rather than through conventional heat flow.
@@
    $\mathbb{U}(1)$-symmetric thermal domain; The right side of the chain is captured by the impurity ansatz with a negative chemical potential, $\mu_{\ell}\leq 0$. The corresponding Wigner functions, illustrated in Fig. 3 (f), display bell-shaped envelopes that can be interpreted as the outcome of a competition between the chemical potential depleting the photon population towards the vacuum, and the thermal fluctuations that sustain a residual population of weakly-interacting photons. Thermodynamically, the energy current is now mostly driven by the strong thermal gradient that has built up in this region of the chain.
Figure 7: RNW regime: destabilization by quantum fluctuations. (a) Transient dynamics of the photon number at the last site of the chain, $n_{L}(t)$, plotted as a function of time $t$ for system sizes $L=100$, $200$, and $400$. The gray line is the classical Gross-Pitaevskii solution for $L=400$. (b) Characteristic timescale $t_{\rm ss}$ to reach the RNW steady state as a function of system size $L$, for three different drive strengths. Each data series is truncated once the RNW regime fails to emerge within the accessible simulation time. The black dashed line is an exponential fit to the data for $F=12.5$. (c) Evidence of metastability near the transition between the RNW and chaotic regimes. The photon number at the last site, $n_{L}(t_{\rm max}\simeq 10^{5})$, is computed from $N_{\rm traj}=200$ Wigner trajectories and time-averaged over a window $\Delta t=10^{3}$. The resulting histogram is shown for $L=200$, $300$, and $400$. In panels (a) and (c), the drive amplitude is fixed at $F=12.5$. The other parameters are set as in Fig. 2.
@@
Remarkably, we argue in Fig. 4 (e) that the sizes of the non-symmetric and thermal domains are limited to the subdominant portions of the chain, while the size of the prethermal domain increases linearly with the total length of the chain $L$. This suggests that the latter may be an extensive thermodynamic phase that occurs ahead of the complete thermalization at the rightmost portion of the chain.
II.4 Regular RNW regime
In this Section, we focus on the RNW regime that was first observed in Ref. [50] in the context of driven-dissipative Klein-Gordon chains coupled to zero-temperature boundary baths, and further studied in Ref. [51]. We discuss how the presence of quantum fluctuations modifies the classical picture.
Performing a quench from the vacuum and evolving under strong coherent drive, the dynamics exhibit two successive temporal regimes, illustrated in Fig. 5 in terms of the average phase $\langle\varphi_{\ell}\rangle$ and photon density $n_{\ell}$. Initially, the system undergoes a prolonged transient during which $\varphi_{\ell}$ is uniformly distributed, rendering $\langle\varphi_{\ell}\rangle$ ill-defined away from the boundary drive. In this stage, $n_{\ell}$ gradually builds up along the chain, starting from the driven boundary site and progressively reaching large values throughout the entire system. This transient corresponds to the chaotic regime discussed in the previous section. It is followed at late times by the onset of the RNW steady state, where $\varphi_{\ell}$ acquires a finite expectation value that slowly rotates along the chain. The dominant wavelength of these oscillations, approximately 30 sites here, as well as the large and uniform photon density, are expected to be dynamically set and independent of the total chain length $L$ [51]. As we show in the Supplementary Information, both $\langle\varphi_{\ell}\rangle$ and $n_{\ell}$ in the RNW steady state closely match their classical counterparts, indicating that these observables do not capture distinct quantum features of the RNW regime. The onset of the RNW steady state is governed by a characteristic timescale $t_{\rm ss}$ that will be discussed below.
II.4.1 Quantum signatures in the steady state
Let us now characterize the distinct quantum signatures of the RNW regime once it is established. In Fig. 6 (a), we monitor the phase fluctuations by means of the circular variance $\Delta\varphi_{\ell}:=1-|\langle{\rm e}^{{\rm i}\varphi_{\ell}}\rangle|$ as a function of the site index $\ell$. In contrast to the driven-dissipative Klein-Gordon chain with zero-temperature boundary baths, where phase fluctuations are absent [50, 51], we find that $\Delta\varphi_{\ell}$ increases monotonically along the chain. This behavior is further illustrated in Fig. 6 (b-e) through the Wigner functions $W_{\ell}\left(\alpha,\alpha^{*}\right)$ at four representative sites in the bulk of the chain; (b) $\ell=5$, (c) $\ell=25$, (d) $\ell=50$, (e) $\ell=95$. The amplitude degree of freedom remains effectively frozen, while the phase undergoes diffusion. As a result, $W_{\ell}\left(\alpha,\alpha^{*}\right)$ is concentrated along an arc in phase space, with its angular extent gradually broadening along the chain. The suppressed fluctuations in the radial direction underpin the sub-Poissonian photon statistics characteristic of the RNW regime.
In Fig. 6 (f), the phase coherence $|g_{1,\ell}^{(1)}|$ exhibits long-range correlations extending across the entire chain. However, contrary to the classical prediction of a uniform $|g_{1,\ell}^{(1)}|$, we observe a slow spatial decay of phase coherence along the chain. This decay arises from quantum fluctuations, which demote the true long-range order predicted in the classical limit. Determining the precise functional form of this decay is challenging due to numerical limitations on accessible system sizes, and because large systems tend to lose the RNW regime, as we discuss below. We also demonstrate that the RNW regime is stable against the presence of weak but finite intrinsic photon losses at all sites, up to decay rates $\gamma_{\rm int}=10^{-3}$. For larger values of $\gamma_{\rm int}$, the RNW first collapses into the chaotic regime and ultimately collapses to a collection of local vacuum states.
II.4.2 Metastability and transition to chaos
We found compelling numerical evidence that the RNW steady state does not persist in the thermodynamic limit in the presence of quantum fluctuations. Following a quench from the vacuum under strong coherent drive, we observe that the emergence of the RNW steady state is governed by a characteristic timescale $t_{\rm ss}$ which increases exponentially with the system size $L$. Moreover, even when evolving the dynamics beyond $t_{\rm ss}$, the system fails to reach the RNW steady state for chains larger than a critical length $L>L^{*}$, instead remaining confined to the chaotic regime described in the previous section.
This is illustrated in Fig. 7 (a), where the photon density at the end of the chain, $n_{L}(t)$, is monitored for various $L$. While $n_{L}(t)$ converges to its RNW value for chains of length $L\leq 200$, it saturates to system-size-dependent values consistent with the chaotic regime for $L\gtrsim 200$. We verified the chaotic nature of this NESS using the diagnostics introduced earlier (data not shown). To highlight the quantum origin of this dynamical phenomenon, we compare with the classical case: the convergence to the RNW steady state is restored when setting the quantum fluctuations to zero in Eqs. (II.1). Moreover, the classical solution exhibits transient chaos before approaching the regular RNW steady state.
Remarkably, within the quantum picture, the size $L^{*}$ at which the breakdown of the RNW regime occurs decreases with increasing drive strength $F$, ruling out a reentrant chaotic phase in the NESS phase diagram as a function of $L$. In Fig. 7 (b), we demonstrate the exponential scaling of $t_{\rm ss}$ that is operationally defined via the convergence criterion $|n_{L}(t>t_{\rm ss})-n_{L}(t_{\rm max})|<\varepsilon$, with $\varepsilon=0.01$, whenever the RNW regime has been reached at the final time $t_{\rm max}$.
We attribute this phenomenon to the emergence of metastability: in the absence of fluctuations, multiple stationary states coexist, namely the chaotic and the RNW steady states [39, 51]. The addition of quantum fluctuations allow trajectories to stochastically transition between these, and the RNW regime becomes metastable for $L>L^{*}$. To probe this, we evolve an ensemble of Wigner trajectories from the vacuum up to a late time $t_{\rm max}$, and subsequently analyze the time-averaged photon number at the chain’s end over a window $\Delta t$. The latter is chosen to suppress high-frequency noise while preserving distinctions between chaotic and regular dynamics. Histograms of this observable, shown in Fig. 7 (c) for a fixed $t_{\rm max}=10^{5}$, $\Delta t=10^{3}$, reveal a clear dynamical crossover when increasing the system size: unimodal at $L=200$ (fully RNW), bimodal at $L=300$ (coexistence), and unimodal (fully chaotic) at $L=400$. Notably, at $L=300$, a significant fraction of the trajectories remains in the chaotic manifold although $t_{\rm max}\gg t_{\rm ss}$.
@@ III Discussion
The NESS of the boundary-driven dissipative chain of quantum nonlinear oscillators revealed an extensive spatial domain, where the chaotic dynamics drive the system to a hydrodynamic regime describable via local equilibria with a large, emergent chemical potential. While consistently with its hydrodynamic nature, this regime displays no distinct quantum signatures, it remarkably exhibits anomalous heating, with temperature increasing away from the coherent drive. More broadly, the emergence of a such a prethermal chaotic phase hinges on three key ingredients: first, an interacting bulk with two conserved charges; second, a local nonequilibrium drive that explicitly breaks the associated symmetries; and finally, weakly-symmetric dissipation channels positioned far from the drive, enabling nontrivial steady states and slow hydrodynamic relaxation. Together with the semiclassical framework developed here, these elements provide a minimal blueprint for realizing prethermal chaotic phases in a broad class of boundary-driven systems.
Quantum fluctuations were found to be relevant to the RNW regime: (i) inducing phase fluctuations along the chain that degrade the long-range coherence of the finite-momentum condensate, and (ii) ultimately destabilizing the condensate in sufficiently long chains. These findings naturally raise the question of how anomalous heating in the prethermal domain and phase fluctuations in the RNW regime impact transport properties, which were reported to be superdiffusive and ballistic, respectively, in classical Klein-Gordon chains [49, 50, 87, 88, 89]. In particular, it is an intriguing possibility that phase fluctuations in the RNW regime might hold fingerprints of Kardar-Parisi-Zhang (KPZ) universality. Also, it remains an open question to assess the role of quantum fluctuations, beyond semiclassics, on both the existence of the prethermal domain and the instability of the RNW regime in the thermodynamic limit. Resolving these challenges will demand either significantly larger numerical simulations or novel approaches beyond TWA, capable of capturing intermediate Kerr nonlinearities while incorporating quantum fluctuations.
@@ IV.1 Stochastic semiclassical description
The TWA is a semiclassical approximation of the quantum many-body dynamics that accounts for leading-order quantum fluctuations. It relies on a phase-space representation of the system’s density matrix $\hat{\rho}$ in terms of the Wigner function $W(\alpha_{1},\alpha_{1}^{*},...,\alpha_{L},\alpha_{L}^{*})$, where $\alpha_{\ell}$ and $\alpha_{\ell}^{*}$ for $\ell=1,...,L$ are the complex amplitudes associated with the local coherent states. In this framework, the Lindblad master equation on the operator $\hat{\rho}$ in Eq. (1) is mapped to a partial differential equation on $W$. Notably, two-body interactions in the Hamiltonian yield contributions up to third order derivatives of the type $\alpha_{\ell}\,\partial^{3}W/\partial\alpha^{*}_{\ell}\partial^{2}\alpha_{\ell}$. The approximation consists of discarding third and higher-order derivatives, reducing the dynamics to a Fokker-Planck equation where $W$ can be interpreted as a probability distribution of the phase-space variables. The latter equation can then be unraveled into a set of $L$ coupled stochastic differential equations on the complex amplitudes $\alpha_{\ell}$ given by Eq. (II.1). In practice, we compute the solutions of these Langevin-like equations, the so-called Wigner trajectories, by means of numerical solvers specific to stochastic differential equations. Observables are computed by averaging over a large number of trajectories generated by different realizations of the quantum noise.
Figure 8: Semiclassical OTOC dynamics. (a) Spatio-temporal scrambling of information in the NESS probed by the semiclassical OTOC $D_{1,\ell}(\tau)$, defined in Eq. (14) and computed via Eq. (6), across a chain of length $L=80$ in the chaotic regime. The dashed line indicates ballistic spreading of information at the butterfly velocity $v=2J$. (b) Dynamics of $D_{1,\ell}(\tau)$ in the chaotic regime computed at sites $\ell=1,10,20,...,80$ (from dark purple to light pink) of an $L=80$ chain. The exponential growth in $\tau$ at a Lyapunov rate $\lambda\simeq 2.8$ is a hallmark of chaotic dynamics. (c) Dynamics of $D_{1,\ell}(\tau)$ in the RNW regime at the same sites as in (b). The sub-exponential growth in time $\tau$ signals non-chaotic dynamics. The dashed line indicates the ergodic bound on the saturation value of the semiclassical OTOC. The results are computed upon averaging over $N_{\rm traj}=5\times 10^{3}$ independent Wigner trajectories. In panels (a) and (b), the drive amplitude is fixed to $F=7.5$, in panel (c) to $F=12.5$. The other parameters are set as in Fig. 2.
The dictionary between the original Lindblad master equation framework and the Wigner-trajectory implementation of the TWA framework reads, notably,
@@
where delimited-⟨$\langle$..$\rangle$ … ⟩ in the observable column is the standard quantum expectation value and, in the Wigner trajectories column, denotes the average with respect to Wigner trajectories. The local Wigner function $W_{\ell}(t;\alpha,\alpha^{*})$ is a phase-space representation of the reduced density matrix at site $\ell$, $\hat{\rho}_{\ell}(t):=\textrm{Tr}_{k\neq\ell}\,\hat{\rho}(t)$. It can be simply reconstructed by generating the histogram of $\alpha_{\ell}(t)$ in the complex plane when sampling over Wigner trajectories. When the NESS is reached, single-time observables converge to constant values and $W_{\ell}(t;\alpha,\alpha^{*})\to W_{\ell}(\alpha,\alpha^{*})$. There, the statistics can be improved by also sampling the trajectories in time, significantly reducing the computational overhead.
IV.2 Semiclassical OTOC
Here we describe in detail our probe of semiclassical chaos, the semiclassical OTOC of Eq. (6). We show how Eq. (6) can be derived from the quantum formulation of the out-of-time order correlators,
$C_{k,\ell}(t,\tau)\!:=\!\frac{1}{2}\textrm{Tr}\Big{(}[\hat{B}_{\ell}(t\!+\!\tau),\hat{A}_{k}(t)]^{\dagger}[\hat{B}_{\ell}(t\!+\!\tau),\hat{A}_{k}(t)]\hat{\rho}(t)\Big{)}.$ (12)
The above OTOC involves a “square commutator” between site $k$ at time $t$ and site $\ell$ at a later time $t+\tau$. The operators $\hat{A}_{\ell}=\hat{n}_{\ell}:=\sum_{n=0}^{\infty}n\,|n\rangle_{\ell}\langle n|_{\ell}$ and $\hat{B}_{\ell}={\rm e}^{{\rm i}\hat{\varphi}_{\ell}}:=\sum_{n=0}^{\infty}|n\rangle_{\ell}\langle n+1|_{\ell}$ are chosen to decompose the local field operator into $\hat{a}_{\ell}=\sqrt{\hat{n}_{\ell}}\,{\rm e}^{{\rm i}\hat{\varphi}_{\ell}}$. They obey the quantum commutation relation $[{\rm e}^{{\rm i}\hat{\varphi}_{k}},\hat{n}_{\ell}]=\delta_{k\ell}\,{\rm e}^{{\rm i}\hat{\varphi}_{\ell}}$. Quantum mechanically, while the absence of a well-defined phase operator is known, the operator ${\rm e}^{{\rm i}\hat{\varphi}_{\ell}}$ is well defined.